OpenStudy (anonymous):

the sum of odd numbers 1-999

OpenStudy (anonymous):

\[\sum_{k=1}^n(2k-1)=n^2\] so your real job is to figure out how many odd numbers you have then square that number

OpenStudy (anonymous):

\[2k-1=999\] \[2k=1000\] \[k=500\] square that one